OpenStudy (anonymous):

inverse of the function y=3x^2, if x> or equal to 0

OpenStudy (bahrom7893):

y=sqrt(x)/3

OpenStudy (campbell_st):

again swap x and y x = 3y^2 x/3 = y^2 \[y = \sqrt{x/3}\] there is a positive and negative root, only consider the positive because of the domain restriction given in the question

OpenStudy (bahrom7893):

woops yea sqrt covers the whole thing